OpenStudy (anonymous):

If you were asked to convert 25 kg to the unit mg, which of the following would be the first fraction used in the conversion?

OpenStudy (disco619):

\[25 \times 1000 \times 1000\] i.e \[25 \times 1000000\]

OpenStudy (disco619):

\[25kg \times 1000 = g\] \[25000g \times 1000 = mg\] \[25kg = 25000000mg\]